If n is any positive odd integer greater than 1, the n(n^2 – 1) is always divisible by:
(a) 7 (b) 5 (c) 24 (d) 15

Answer by richwmiller(17219)

n(n^2 – 1)
3*(9-1)=24
5*(25-1)=5*24
7*(49-1)=7*48
You can decide now